# Interplay of Zeeman field, Rashba spin-orbit interaction, and superconductivity: Transition temperature and quasiparticle excitations

### v1

@@ I Introduction
As an example, the Knight shift and the spin-lattice relaxation rate $1/T_{1}$ in nuclear magnetic resonance (NMR) are crucial to assessing spin susceptibility in superconducting states, which in turn reveals the pairing symmetry through spin channels. The effectiveness of NMR is illustrated below. (i) When the temperature decreases, the Knight shift decreases below $T_{c}$ and stays isotropic, eventually vanishing at zero temperature for an $s$-wave and spin-singlet pairing state [2]. Conversely, for an $p$-wave and spin-triplet pairing state, it shows anisotropic temperature dependence, decreasing when the external magnetic field is oriented in one direction, while remaining unchanged in another direction [3]. (ii) The spin-lattice relaxation rate $1/T_{1}$ shows a Hebel-Slichter coherence peak at $T=T_{c}$ in an $s$-wave pairing state with a pairing function that remains sign-unchanged on the Fermi surface. This coherence peak is diminished or absent in a sign-changing pairing state, such as a $p$-wave pairing state [3, 4, 5, 6].
In the majority of theoretical investigations, an electric or magnetic field imposed on a quantum many-particle system is considered as a perturbation, resulting in linear response theory. However, in NMR measurements, a substantial signal is detectable only with a strong external magnetic field. The energy scale of such a magnetic field can match or surpass the superconducting energy gap; thus, the applied magnetic field should not be considered a mere perturbation. Instead, one must determine the ground state of the many-particle system under the influence of this external magnetic field. Consequently, the system’s responses to the external magnetic field should be assessed based on this revised ground state.
@@
The spin-orbit coupling (SOC) of electrons arises from relativistic effects, as elaborated in Ref. [9]. The corresponding Hamiltonian is given by
$H_{\text{SO}}=\frac{1}{2m^{2}c^{2}}\left(\nabla{}V\times{}\mathbf{p}\right)\cdot\mathbf{S},$
where $V$ represents the periodic potential experienced by the electron, $\mathbf{p}$ is its momentum, and $\mathbf{S}$ indicates its spin. In the context of symmetry-based energy band theory, notably through $\mathbf{k}\cdot\mathbf{p}$ perturbation theory, the spin-orbit interaction for an individual electron is conventionally represented as:
$H_{\text{SO}}=\mathbf{g}_{\mathbf{k}}\cdot\hat{\sigma},$
where $\hat{\sigma}$ is a vector consisting of the three Pauli matrices, and $\mathbf{g}_{\mathbf{k}}$ includes only odd powers of the linear momentum $\mathbf{k}$. Furthermore, this paper examines a simplified form of SOC, defined by
@@
with the particular simplification $\mathbf{g}_{\mathbf{k}}=g\mathbf{k}$. This reduction is sufficient to demonstrate substantial SOC effects in a superconductor.
@@
$H_{P}$ Pauli limit of $s$-wave pairing state, $H_{P}=H_{0}=\Delta_{0}/\sqrt{2}\mu_{B}$
@@ II Theoretical model and formulation
Normal state: Consider the normal state of the system, which is governed by the unpaired part $H_{0}$. In the presence of an external Zeeman field $\mathbf{H}$ and Rashba SOC with coupling strength $g$, the unpaired part $H_{0}$ is given by
@@
$\lvert\mathbf{k},+\rangle=\lvert\mathbf{k}\rangle\otimes\begin{pmatrix}\cos\frac{\Theta_{\mathbf{k}}}{2}\\ \sin\frac{\Theta_{\mathbf{k}}}{2}e^{i\Phi_{\mathbf{k}}}\end{pmatrix},\,\lvert\mathbf{k},-\rangle=\lvert\mathbf{k}\rangle\otimes\begin{pmatrix}\sin\frac{\Theta_{\mathbf{k}}}{2}\\ -\cos\frac{\Theta_{\mathbf{k}}}{2}e^{i\Phi_{\mathbf{k}}}\end{pmatrix}.$ (5b)
@@
$\begin{split}\Delta_{++}=&\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+e^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right)+e^{-i\Phi_{\mathbf{k}}}\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+e^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right),\\ \Delta_{--}=&\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-e^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right)-e^{-i\Phi_{\mathbf{k}}}\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-e^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right),\\ \Delta_{+-}=&\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-e^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right)+e^{-i\Phi_{\mathbf{k}}}\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-e^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right),\\ \Delta_{-+}=&\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+e^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right)-e^{-i\Phi_{\mathbf{k}}}\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+e^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right).\end{split}$ (14)
@@ III $s$-wave pairing state
in the pairing potential as described in Eq. (9). For the $s$-wave pairing function specified in Eq. (11), the gap equation (7) simplifies to the form
$F_{s}-F_{n}=-\frac{1}{\beta}\sum_{\mathbf{k}}\ln\frac{(1+e^{-\beta E_{\mathbf{k}+}})(1+e^{-\beta E_{\mathbf{k}-}})}{(1+e^{-\beta\xi_{\mathbf{k}+}})(1+e^{-\beta\xi_{\mathbf{k}-}})}+\frac{\lvert\Delta\lvert^{2}}{V_{0}}+\sum_{\mathbf{k}}\left[\xi_{\mathbf{k}}-\frac{E_{-\mathbf{k}+}+E_{-\mathbf{k}-}}{2}\right],$ (16)
$F_{s}(T=T_{c})-F_{n}(T=T_{c})=0,$ (17)
where $F_{s}-F_{n}$ is specified in Eq. (16).
@@ III.1 Superconducting transition temperature
Zeeman field effect: Assuming without loss of generality, we consider a Zeeman field $\mathbf{H}=H_{z}\hat{z}$. The Zeeman field aligns the electron spin, which hinders the formation of a spin-singlet Cooper pair and reduces $T_{c}$. When $H_{z}=0$, the superconducting $T_{c}$ reaches its maximum value of $T_{c0}=T_{c}(H_{z}=0)$, but when the Zeeman field $H_{z}$ exceeds the Pauli limit [29, 30], the superconducting $T_{c}$ vanishes. According to Maki and Tsuneto [30], there exists a critical threshold of $T_{0}=0.556T_{c0}$, which dictates whether the superconductor to normal state phase transition is second or first order. Specifically, the phase transition is of second order when $T_{c}>T_{0}$, and becomes of first order when $T_{c}<T_{0}$.
@@
$\Delta(\mathbf{k})=\left\{\begin{aligned} &\,-\frac{1}{2}\Delta e^{-i\varphi_{\mathbf{k}}}(\sigma_{0}+\sigma_{z}),&\mathbf{k}\in I_{1},\\ &\,-\Delta e^{-i\varphi_{\mathbf{k}}}\sigma_{z},&\mathbf{k}\in I_{2},\\ &\,\frac{1}{2}\Delta e^{-i\varphi_{\mathbf{k}}}(\sigma_{0}-\sigma_{z}),&\mathbf{k}\in I_{3}.\end{aligned}\right.$ (19)
This diagonal form without interband pairing is anticipated in the basis formed by the Kramers pair and is owing to the rotational symmetry of the $s$-wave pairing. Furthermore, we can conclude that the diagonal form of the pairing function $\Delta(\mathbf{k})$ results in a gap equation independent of the strength of Rashba SOC (see Appendix C for more details). Therefore, both the gap function $\Delta(T)$ and the superconducting $T_{c}$ remain unchanged, as confirmed by numerical verification.
@@
$\begin{split}&\Delta_{++}=-\Delta_{--}=-\Delta{}e^{-i\varphi_{\mathbf{k}}}\sin{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}},\\ &\Delta_{+-}=\Delta_{-+}=\Delta{}e^{-i\varphi_{\mathbf{k}}}\cos{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}}.\end{split}$ (20)
@@ III.2 Quasiparticles: Bogoliubov Fermi surface
The energy spectrum of Bogoliubov quasiparticles can be found by diagonalizing the $4\times 4$ matrix (see Appendix C). The explicit form can be derived by specifying the polar angle $\theta_{\mathbf{k}}=0,\pi/2,\pi$:
@@ III.3 Finite total momentum pairing: FFLO state
In the end of this section, we would like to explore the feasibility of a finite total momentum Cooper pairing state, such as the prominent Fulde-Ferrell (FF) [31] and Larkin-Ovchinnikov (LO) [32] states, together referred to as FFLO, or pair density waves as discussed in recent studies (see Ref. [33] for a comprehensive review). Specifically, for an $s$-wave pairing superconducting state where each Cooper pair has a total momentum $\mathbf{q}$, one can derive a similar self-consistent gap equation for the pairing function $\Delta_{\mathbf{q}}$, solve it and compute the condensation energy $E_{c}$, defined as the energy difference between the normal and superconducting ground states, as a function of $q$. Thus, an FFLO ground state is identified when the condensation energy $E_{c}$ reaches a maximum at a finite $q=|\mathbf{q}|$. The technical details can be found in Appendix D
@@ IV $p$-wave pairing states
Table 2: Examples of $p$-wave pairing states. Here $Y_{1,\pm 1}(\hat{\mathbf{k}})=\sqrt{\frac{3}{8\pi}}\sin{\theta_{\mathbf{k}}}e^{\pm i\varphi_{\mathbf{k}}}$ and $Y_{1,0}(\hat{\mathbf{k}})=\sqrt{\frac{3}{4\pi}}\cos{\theta_{\mathbf{k}}}$ are spherical harmonics.
@@
$\mathbf{d}(\mathbf{k})$ $\Delta\sin\theta_{\mathbf{k}}e^{i\varphi_{\mathbf{k}}}\hat{z}$ $\Delta\cos\theta_{\mathbf{k}}\hat{z}$ $\Delta\sin\theta_{\mathbf{k}}\left(\cos\phi_{\mathbf{k}}\hat{x}+\sin\phi_{\mathbf{k}}\hat{y}\right)$
@@
The $p$-wave and spin-triplet pairing states are classified into opposite-spin pairing (OSP) and equal-spin pairing (ESP) categories [3]. Table 2 provides examples of these $p$-wave pairing states. It is important to note that this classification depends on the chosen spin quantization axis and can change as a result. For instance, consider the following relation:
@@
where $\left|\uparrow_{\alpha}\right\rangle$ and $\left|\downarrow_{\alpha}\right\rangle$ are defined as $\sigma_{\alpha}\left|\uparrow_{\alpha}\right\rangle=+\left|\uparrow_{\alpha}\right\rangle$ and $\sigma_{\alpha}\left|\downarrow_{\alpha}\right\rangle=-\left|\downarrow_{\alpha}\right\rangle$ respectively, with $\alpha=x,z$. (For further information, refer to Appendix E.) To discuss OSP and ESP states, the spin quantization axis must be established. Throughout this paper, unless specified differently, we will align the spin quantization axis along the $z$-axis, especially when there is no external Zeeman field, or along the direction of the imposed Zeeman field.
@@
Quasiparticle energy dispersion: In the presence of both the Zeeman field and Rashba SOC, deriving a general analytical expression for the quasiparticle energy dispersion is challenging. However, when either the Zeeman field or Rashba SOC is present alone, such an expression can be obtained.

### v2

@@ I Introduction
As an example, the Knight shift and the spin-lattice relaxation rate $1/T_{1}$ in nuclear magnetic resonance (NMR) are crucial to assessing spin susceptibility in superconducting states, which in turn reveals the pairing symmetry through spin channels [2, 3, 4, 5, 6]. In the majority of theoretical investigations, an electric or magnetic field imposed on a quantum many-particle system is considered as a perturbation, resulting in linear response theory. However, in NMR measurements, a substantial signal is detectable only with a strong external magnetic field. The energy scale of such a magnetic field can match or surpass the superconducting energy gap; thus, the applied magnetic field should not be considered a mere perturbation. Instead, one must determine the ground state of the many-particle system under the influence of this external magnetic field. Consequently, the system’s responses to the external magnetic field should be assessed based on this revised ground state.
@@
The spin-orbit coupling (SOC) of electrons arises from relativistic effects, as elaborated in Ref. [9]. In the context of symmetry-based energy band theory, notably through $\mathbf{k}\cdot\mathbf{p}$ perturbation theory, the spin-orbit interaction for an individual electron is conventionally represented as: $H_{\text{SO}}=\mathbf{g}_{\mathbf{k}}\cdot\hat{\sigma}$, where $\hat{\sigma}$ is a vector consisting of the three Pauli matrices, and $\mathbf{g}_{\mathbf{k}}$ includes only odd powers of the linear momentum $\mathbf{k}$. Furthermore, this paper examines a simplified form of SOC, defined by
@@
with the particular simplification $\mathbf{g}_{\mathbf{k}}=g\mathbf{k}$. It is noted that this form is straightforward, preserving time reversal symmetry while breaking both spin rotational and spatial inversion symmetry. Although it does not exactly align with the specifics of a typical non-centrosymmetric superconductor, as a leading term, it aptly captures the fundamental physics and is commonly used in theoretical studies [10, 11, 12, 13].
@@
$H_{P}$ Pauli limit of $s$-wave pairing state, $H_{P}\approx H_{0}=\Delta_{0}/\sqrt{2}\mu_{B}$
@@ II Theoretical model and formulation
Normal state: First, we shall consider the normal state of the system, which is governed by the unpaired part $H_{0}$. In the presence of an external Zeeman field $\mathbf{H}$ and Rashba SOC with coupling strength $g$, the unpaired part $H_{0}$ is given by
@@
$\lvert\mathbf{k},+\rangle=\lvert\mathbf{k}\rangle\otimes\begin{pmatrix}\cos\frac{\Theta_{\mathbf{k}}}{2}\\ \sin\frac{\Theta_{\mathbf{k}}}{2}\mathrm{e}^{i\Phi_{\mathbf{k}}}\end{pmatrix},\,\lvert\mathbf{k},-\rangle=\lvert\mathbf{k}\rangle\otimes\begin{pmatrix}\sin\frac{\Theta_{\mathbf{k}}}{2}\\ -\cos\frac{\Theta_{\mathbf{k}}}{2}\mathrm{e}^{i\Phi_{\mathbf{k}}}\end{pmatrix}.$ (5b)
@@
$\begin{split}\Delta_{++}=&\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right)+\mathrm{e}^{-i\Phi_{\mathbf{k}}}\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right),\\ \Delta_{--}=&\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right)-\mathrm{e}^{-i\Phi_{\mathbf{k}}}\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right),\\ \Delta_{+-}=&\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right)+\mathrm{e}^{-i\Phi_{\mathbf{k}}}\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}-\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}\right),\\ \Delta_{-+}=&\sin\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\uparrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\uparrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right)-\mathrm{e}^{-i\Phi_{\mathbf{k}}}\cos\frac{\Theta_{\mathbf{k}}}{2}\left(\Delta_{\downarrow\uparrow}\cos\frac{\Theta_{-\mathbf{k}}}{2}+\mathrm{e}^{-i\Phi_{-\mathbf{k}}}\Delta_{\downarrow\downarrow}\sin\frac{\Theta_{-\mathbf{k}}}{2}\right).\end{split}$ (14)
@@ III $s$-wave pairing state
in the pairing potential as described in Eq. (9).
@@ III.1 Superconducting transition temperature
Zeeman field effect: Assuming without loss of generality, we consider a Zeeman field $\mathbf{H}=H_{z}\hat{z}$. The Zeeman field aligns the electron spin, which hinders the formation of a spin-singlet Cooper pair and reduces $T_{c}$. When $H_{z}=0$, the superconducting $T_{c}$ reaches its maximum value of $T_{c0}=T_{c}(H_{z}=0)$, but when the Zeeman field $H_{z}$ exceeds the Pauli limit [29, 30], the superconducting $T_{c}$ vanishes. According to Maki and Tsuneto [30], there exists a critical threshold of $T_{0}=0.556T_{c0}$, which dictates whether the superconductor to normal state phase transition is second or first order. Specifically, the phase transition is second order when $T_{c}>T_{0}$, and becomes of first order when $T_{c}<T_{0}$.
@@
$\Delta(\mathbf{k})=\left\{\begin{aligned} &\,-\frac{1}{2}\Delta\mathrm{e}^{-i\varphi_{\mathbf{k}}}(\sigma_{0}+\sigma_{z}),&\mathbf{k}\in I_{1},\\ &\,-\Delta\mathrm{e}^{-i\varphi_{\mathbf{k}}}\sigma_{z},&\mathbf{k}\in I_{2},\\ &\,\frac{1}{2}\Delta\mathrm{e}^{-i\varphi_{\mathbf{k}}}(\sigma_{0}-\sigma_{z}),&\mathbf{k}\in I_{3}.\end{aligned}\right.$ (16)
This diagonal form without interband pairing is anticipated in the basis formed by the Kramers pair and is owing to the rotational symmetry of the $s$-wave pairing. Furthermore, we can conclude that the diagonal form of the pairing function $\Delta(\mathbf{k})$ results in a gap equation independent of the strength of Rashba SOC (see Appendix D for more details). Therefore, both the gap function $\Delta(T)$ and the superconducting $T_{c}$ remain unchanged, as confirmed by numerical verification.
@@
$\begin{split}&\Delta_{++}=-\Delta_{--}=-\Delta{}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\sin{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}},\\ &\Delta_{+-}=\Delta_{-+}=\Delta{}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\cos{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}}.\end{split}$ (17)
@@ III.2 Quasiparticles: Bogoliubov Fermi surface
The energy spectrum of Bogoliubov quasiparticles can be found by diagonalizing the $4\times 4$ matrix (see Appendix D). The explicit form can be derived by specifying the polar angle $\theta_{\mathbf{k}}=0,\pi/2,\pi$:
@@ III.3 Finite total momentum pairing: FFLO state
In the end of this section, we would like to explore the feasibility of a finite total momentum Cooper pairing state, such as the prominent Fulde-Ferrell (FF) [31] and Larkin-Ovchinnikov (LO) [32] states, together referred to as FFLO, or pair density waves as discussed in recent studies (see Ref. [33] for a comprehensive review). Specifically, for an $s$-wave pairing superconducting state where each Cooper pair has a total momentum $\mathbf{q}$, one can derive a similar self-consistent gap equation for the pairing function $\Delta_{\mathbf{q}}$, solve it and compute the condensation energy $E_{c}$, defined as the energy difference between the normal and superconducting ground states, as a function of $q$. Thus, an FFLO ground state is identified when the condensation energy $E_{c}$ reaches a maximum at a finite $q=|\mathbf{q}|$. The technical details can be found in Appendix E
@@ IV $p$-wave pairing states
Table 2: Examples of $p$-wave pairing states. Here $Y_{1,\pm 1}(\hat{\mathbf{k}})=\sqrt{\frac{3}{8\pi}}\sin{\theta_{\mathbf{k}}}\mathrm{e}^{\pm i\varphi_{\mathbf{k}}}$ and $Y_{1,0}(\hat{\mathbf{k}})=\sqrt{\frac{3}{4\pi}}\cos{\theta_{\mathbf{k}}}$ are spherical harmonics.
@@
$\mathbf{d}(\mathbf{k})$ $\Delta\sin\theta_{\mathbf{k}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\hat{z}$ $\Delta\cos\theta_{\mathbf{k}}\hat{z}$ $\Delta\sin\theta_{\mathbf{k}}\left(\cos\phi_{\mathbf{k}}\hat{x}+\sin\phi_{\mathbf{k}}\hat{y}\right)$
@@
The $p$-wave and spin-triplet pairing states are classified into opposite-spin pairing (OSP) and equal-spin pairing (ESP) categories [3]. Table 2 provides examples of these $p$-wave pairing states. It is important to note that this classification depends on the chosen spin quantization axis and can change as a result. For instance, by taking the spin quantization axis along z and x axes, we obtain the following relation:
@@
where $\left|\uparrow_{\alpha}\right\rangle$ and $\left|\downarrow_{\alpha}\right\rangle$ are defined as $\sigma_{\alpha}\left|\uparrow_{\alpha}\right\rangle=+\left|\uparrow_{\alpha}\right\rangle$ and $\sigma_{\alpha}\left|\downarrow_{\alpha}\right\rangle=-\left|\downarrow_{\alpha}\right\rangle$ respectively, with $\alpha=x,z$. (For further information, refer to Appendix F.) To discuss OSP and ESP states, the spin quantization axis must be established. Throughout this paper, unless specified differently, we will align the spin quantization axis along the $z$-axis, especially when there is no external Zeeman field, or along the direction of the imposed Zeeman field.
@@
In the presence of both the Zeeman field and Rashba SOC, deriving a general analytical expression for the quasiparticle energy dispersion is challenging. However, when either the Zeeman field or Rashba SOC is present alone, such an expression can be obtained.